The length of a rectangle is 4 more than its width, and the ratio of the length to its
width is 5 to 3. Find the dimensions of the rectangle.

Found 2 solutions by dkppathak, FrankM:
Answer by dkppathak(439)

length is 4 more than width
L=4+B i
L/B=5/3 ii
B =3L/5 by putting value in equation i
L =4+3l/5
L-3L/5 =4
2L/5 =4
2 L =5x4 =20
L=20/2
L=10
B= L-4 =10-4 = 6

ANSWER L=10 cm
B=6 cm

Answer by FrankM(1040)

L=W+4
L/W = 5/3
Substitue for L
(W+4)/W = 5/3
Cross multiply
5W=3(W+4)
5W=3W+12
2W=12
W=6, L=10